OpenStudy (anonymous):

please help!!! if i have this sequence 3,6,9,12..increasing by 3..what would be the arithmetic sequence, explicit sequence and recursive formula..HELP!

OpenStudy (anonymous):

i think the explicit formula would be an=a1+(n-1)(d) an=3+(n-1)(3) an=20+3n-3 an=3n+17 ...but i'm not sure

OpenStudy (anonymous):

@hartnn do you know how to do this? emergency!!

hartnn (hartnn):

an=3+(n-1)(3) then what u did ?

OpenStudy (anonymous):

lol i'm not sure, i did that a few days ago..uh, i have no idea where i got the 20 from:S

hartnn (hartnn):

an=3+(n-1)(3) = 3+ 3n -3 =.... ?

OpenStudy (anonymous):

en?

OpenStudy (anonymous):

3n

hartnn (hartnn):

es! thats your explicit formula :)

OpenStudy (anonymous):

yay, so how would i do recursive formula?

OpenStudy (anonymous):

i think the formula is an=an-1+d but i'm not sure

hartnn (hartnn):

hmm...recursive is even simpler. 'each term increasing by 3' means if you add 3 to current, you get next term, did u get this ?

OpenStudy (anonymous):

yess

hartnn (hartnn):

an=an-1+d is correct formula, where you just need value of d

hartnn (hartnn):

\(\large a_n=a_{n-1}+d\)

hartnn (hartnn):

d=... ?

OpenStudy (anonymous):

3

hartnn (hartnn):

yes! so recursive formula is just \(a_n=a_{n-1}+3\) thats it!

OpenStudy (anonymous):

seriously? woahh :) thanks! and the last one is arithmetic sequence?

hartnn (hartnn):

last one ?

OpenStudy (anonymous):

yess, i just needed to know how to do them three :)

hartnn (hartnn):

Question : what would be the arithmetic sequence, explicit sequence and recursive formula Answer : arithmetic sequence: 3,6,9,12.. (given) explicit formula : an = 3n recursive formula : an = a(n-1)+3 what else is remaining ?

OpenStudy (anonymous):

uh, the a?

OpenStudy (precal):

this is also recursive